A 1210-turn solenoid 29.2 cm long and 1.58 cm in diameter carries 165 mA.

How much magnetic energy does it contain?

Express your answer with the appropriate units.
Answer #1

energy in a solenoid = 1/2* L *I^2

L = u0 * n^2 * A/l = 1.2354 mH

so

energy = 1/2 * 1.2354 *10^-3 * 165 *10^-3

                 = 101.92 * 10^-6 J